OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y - z = 2 3x + y - 2z = -15 3x + 4y - 3z = -2

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

hero help me please

OpenStudy (anonymous):

you don't have't to use substitute method i just need the answer

OpenStudy (anonymous):

it is easy so do it : x=-2y+z+2 substitute this in the rest of two equations , after that , substitute the y and so on

OpenStudy (anonymous):

can you just give me the answer, or show me your work please

OpenStudy (anonymous):

roh please give me the answer, and i need some help with other ones

OpenStudy (anonymous):

ok wait i am solving just need some more time

OpenStudy (anonymous):

ok thank, i'll post another question on here

OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. 2x + 4y - 3z = -7 3x + y + 4z = -12 x + 3y + 4z = 4

OpenStudy (anonymous):

i'll fan you too

OpenStudy (anonymous):

ok i got y=+9 z=+24 searching for x

OpenStudy (anonymous):

a. (7, -4, -1) b. (-7, -4, 1) c. (7, -4, 1) d. (-7, 4, -1) theses are the possible answers

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

i think yor doing it wrong

OpenStudy (anonymous):

may be let me try once again

OpenStudy (anonymous):

ya, those are the answers up there it could be,

OpenStudy (anonymous):

i jsut need the answer to the first and second question i asked thats it,

OpenStudy (anonymous):

x=-7, y=4, z=-1

OpenStudy (anonymous):

i 100% sure option D is the answer